# Guided Diffusion Sampling on Function Spaces with Applications to PDEs

### v1

@@ Abstract
We propose a general framework for conditional sampling in PDE-based inverse problems, targeting the recovery of whole solutions from extremely sparse or noisy measurements. This is accomplished by a function-space diffusion model and plug-and-play guidance for conditioning. Our method first trains an unconditional discretization-agnostic denoising model using neural operator architectures. At inference, we refine the samples to satisfy sparse observation data via a gradient-based guidance mechanism. Through rigorous mathematical analysis, we extend Tweedie’s formula to infinite-dimensional Banach spaces, providing the theoretical foundation for our posterior sampling approach. Our method (FunDPS) accurately captures posterior distribution in function spaces under minimal supervision and severe data scarcity. Across five PDE tasks with only 3% observation, our method achieves an average 32% accuracy improvement over state-of-the-art fixed-resolution diffusion baselines while reducing sampling steps by 4x. Furthermore, multi-resolution fine-tuning ensures strong cross-resolution generalizability. To the best of our knowledge, this is the first diffusion-based framework to operate independently of discretization, offering a practical and flexible solution for forward and inverse problems in the context of PDEs. Code is available at https://github.com/neuraloperator/FunDPS.
@@ 1 Introduction
2) Mixed-resolution conditional generation pipeline in function space (FunDPS).
We propose the function-space diffusion posterior sampling method that builds on the above theoretical framework, the pipeline of which is shown in Figure 2. An unconditional diffusion model is first pretrained using the score matching objective as defined in [14]. The model works on a joint representation of PDE parameters and solution, which allows us to solve forward problems, inverse problems, and a combination thereof using partial observations. The model begins training on a coarse grid, then fine-tunes on a finer one. The Gaussian random field (GRF)-based noise model and neural operator architecture enable this process with minimal accuracy loss. After training the unconditional model, we generate conditional samples as follows: The sample is initialized with GRF noise, followed by iterative denoising via a second-order deterministic sampler [16]. At each step, the sample evolves according to the conditional score operator, which decomposes into data prior (from the pretrained model) and observation-based likelihoods. Our generalized Tweedie’s formula enables us to approximate the gradient of the observation function w.r.t the current sample for log-likelihood calculation. FunDPS can also employ mixed-resolution inference by performing most sampling steps at a lower resolution and only upsampling towards the end, all while maintaining accuracy.
@@
We test our approach on five challenging PDEs, which vary widely in their difficulty due to complex input distributions (GRF’s multi-scale features), different boundary conditions (Dirichlet and periodic), and nonlinear patterns (Navier-Stokes). We simulate extreme obfuscation in measurements by revealing only 3% of the function’s spatial domain. Specifically, for forward tasks (e.g., recovering the final state), we only observe $3\%$ of the initial condition and reconstruct the entire solution, while for inverse tasks, we observe only $3\%$ of the final state and reconstruct unknown initial or coefficient fields. Despite these difficulties, our method consistently achieves top performance (see Figure 1). Across all five tasks, we reduce the average error by roughly 32% compared to the state-of-the-art diffusion-based solver (DiffusionPDE [17], which operates at a fixed trained resolution) and surpass deterministic neural PDE solvers by even larger margins. Moreover, our reverse diffusion process requires only 200–500 steps to converge—up to 10$\times$ fewer than DiffusionPDE—while maintaining superior accuracy (see Figure 4a). We attribute these improvements primarily to FunDPS’s function-space formulation, which learns a more expressive prior and applies smoother guidance, thereby moving samples toward the posterior in significantly fewer steps. Additionally, thanks to its neural-operator architecture, FunDPS utilizes multi-resolution training and inference pipelines, which reduces GPU training hours by $\sim$25compared to full-resolution pretraining, and cuts inference wall-clock time by 25$\times$ compared to DiffusionPDE with minimal impact on accuracy.
@@ 3.1 Problem Settings
We assume that the measurement noise $\varepsilon$ is a Gaussian random field (GRF) with covariance operator $\mathbf{C}_{\eta}$, i.e., $\varepsilon$ is drawn from the Gaussian measure $\eta=\mathcal{N}(0,\mathbf{C}_{\eta})$ on $\mathcal{U}$, which is the natural extension of Gaussian noise to function spaces [50]. Observe that, given ${\boldsymbol{a}}$, we have that ${\boldsymbol{u}}\mid{\boldsymbol{a}}$ ${}\sim\eta^{{\boldsymbol{A}}({\boldsymbol{a}})}=\mathcal{N}({\boldsymbol{A}}({\boldsymbol{a}}),\mathbf{C}_{\eta})$ is drawn from a Gaussian measure $\eta^{{\boldsymbol{A}}({\boldsymbol{a}})}$ with mean ${\boldsymbol{A}}({\boldsymbol{a}})$. When the noise-free observation ${\boldsymbol{A}}({\boldsymbol{a}})$ is an element of the Cameron-Martin space $\mathbf{C}_{\eta}^{1/2}(\mathcal{U})=:H(\eta)\subset\mathcal{U}$ associated with $\mathbf{C}_{\eta}$, Cameron-Martin theorem allows us to compute the Radon-Nikodym derivative 222 We recall that the inner product and norm on the Cameron-Martin space $\mathcal{U}_{0}$ are readily computed via $\langle{\boldsymbol{u}}_{0},{\boldsymbol{u}}_{1}\rangle_{H(\eta)}=\langle\mathbf{C}_{\eta}^{-1/2}{\boldsymbol{u}}_{0},\mathbf{C}_{\eta}^{-1/2}{\boldsymbol{u}}_{1}\rangle_{\mathcal{U}},\qquad{\boldsymbol{u}}_{0},{\boldsymbol{u}}_{1}\in H(\eta).$
$\frac{\,\mathrm{d}\eta^{{\boldsymbol{A}}({\boldsymbol{a}})}}{\,\mathrm{d}\eta}({\boldsymbol{u}})=\exp\left(\langle{\boldsymbol{u}},{\boldsymbol{A}}({\boldsymbol{a}})\rangle_{H(\eta)}-\tfrac{1}{2}\lVert{{\boldsymbol{A}}({\boldsymbol{a}})}\rVert^{2}_{H(\eta)}\right).$ (3)
In other words, Eq. 3 gives the density of $\eta^{\boldsymbol{a}}$ with respect to the noise measure $\eta$. We will use $\Phi:\mathcal{A}\times\mathcal{U}\to\mathbb{R}$ to represent the function $\Phi({\boldsymbol{a}},{\boldsymbol{u}})=\langle{\boldsymbol{u}},{\boldsymbol{A}}({\boldsymbol{a}})\rangle_{\mathcal{U}_{0}}-\tfrac{1}{2}\lVert{\boldsymbol{A}}({\boldsymbol{a}})\rVert^{2}_{\mathcal{U}_{0}}$. Loosely speaking, $\Phi({\boldsymbol{a}},{\boldsymbol{u}})$ is the log-likelihood of ${\boldsymbol{u}}$ given ${\boldsymbol{a}}$. In the special case that ${\boldsymbol{u}}\in\mathbb{R}^{n}$ is finite dimensional (for instance, when ${\boldsymbol{A}}$ is the composition of a PDE solution operator and a finite observation mask), the measure $\eta$ corresponds to a mean-zero Gaussian random variable with covariance matrix $\mathbf{C}_{\eta}\in\mathbb{R}^{n\times n}$ In this case, when $\mathbf{C}_{\eta}$ has full rank, the Cameron-Martin space is $\mathbf{C}_{\eta}^{1/2}(\mathbb{R}^{n})=\mathbb{R}^{n}$ so that the log-likelihood is defined for any value of ${\boldsymbol{A}}({\boldsymbol{a}})$.
@@
We propose to learn the prior distribution $\nu({\boldsymbol{a}})$ from data using the recent extension of diffusion models to function spaces [14, 51]. These methods define a sequence of distributions $\nu_{t}({\boldsymbol{a}}_{t})$ that progressively add noise to the data until approximately reaching a tractable latent distribution $\mu_{T}\approx\mathbf{\Gamma}$. Learning a score approximator ${\boldsymbol{D}}_{\theta}({\boldsymbol{a}}_{t},t)\approx\mathbb{E}[{\boldsymbol{a}}_{0}|{\boldsymbol{a}}_{t}]$ from the data via a variant of score matching, one can approximately reverse the noising process [^3] Following common conventions, we will also refer to the noising process as forward process. and sample from the data distribution. Moreover, by replacing the score with a conditional expectation, one can solve the reverse SDE given initial conditions [43]. Notably, the denoiser ${\boldsymbol{D}}_{\theta}$ is parametrized as a neural operator [14, 44, 52], and the noise is chosen to be Gaussian random fields, as opposed to neural networks and multivariate Gaussian random variables for the finite-dimensional setting (see Appendix A for more details).
@@ 3.2 Conditional Score via Likelihood and Prior Score
for some specified time-dependent constants $\sigma_{t}>0$ [14]. Here, $\gamma=\mathcal{N}(0,\mathbf{C}_{\gamma})$ is now a Gaussian measure on the space $\mathcal{A}$. We will also use $\gamma_{t}=\mathcal{N}(0,\sigma_{t}^{2}\mathbf{C}_{\gamma})$ to represent the corresponding Gaussian measure with scaled covariance. The spaces $H(\gamma),H(\gamma_{t})$ represent the Cameron-Martin spaces associated with these Gaussian measures. We will write $\nu_{0}({\boldsymbol{a}}_{0})=\nu({\boldsymbol{a}})$ for the prior distribution over ${\boldsymbol{a}}$ and $\nu_{t}({\boldsymbol{a}}_{t})$ for the marginal distribution of ${\boldsymbol{a}}_{t}$ obtained during the forward process. Similarly, the conditional forward process is obtained in the same manner as Eq. 4 except with initial conditions ${\boldsymbol{a}}_{0}\sim\nu^{{\boldsymbol{u}}}$ drawn from the posterior corresponding to a given, fixed ${\boldsymbol{u}}$. We will use $\nu_{0}^{{\boldsymbol{u}}}=\nu^{{\boldsymbol{u}}}$ and $\nu_{t}^{{\boldsymbol{u}}}$ for the corresponding measures.
As in prior work [14, 43], we make the assumption that the prior $\nu_{0}$ is supported on $H(\gamma)$, i.e., $\nu(H(\gamma))=1$. Note that as the posterior $\nu_{0}^{{\boldsymbol{u}}}\ll\nu_{0}$ is absolutely continuous with respect to the prior, we also have that $\nu_{0}^{{\boldsymbol{u}}}(H(\gamma))=1$. Under this assumption, both $\nu_{t}$ and $\nu_{t}^{{\boldsymbol{u}}}$ are equivalent to $\gamma_{t}$ in the sense of mutual absolute continuity (see Appendix B and [14, Lemma 13]).
@@ 3.3 Approximating the Likelihood in Function Spaces
While the calculations in the previous section in principle allow for plug-and-play posterior sampling, a major difficulty is that the time-dependent likelihood $\tilde{\Phi}_{t}$ is intractable. In particular, from Eq. 3 we know the log-likelihood of ${\boldsymbol{u}}$ given a noise-free sample ${\boldsymbol{a}}_{0}$, but Eq. 6 requires us to have access to the log-likelihood of ${\boldsymbol{u}}$ given a noisy sample ${\boldsymbol{a}}_{t}$. In the next two sections, we develop a method for approximating this likelihood through a conditional expectation.
Let us write $\mu_{t}^{{\boldsymbol{a}}_{t}}$ for the conditional distribution of ${\boldsymbol{u}}\mid{\boldsymbol{a}}_{t}$. This measure can be sampled from by firt predicting the clean ${\boldsymbol{a}}_{0}$ from the noisy ${\boldsymbol{a}}_{t}$, followed by sampling from the noise measure $\eta^{{\boldsymbol{A}}({\boldsymbol{a}}_{0})}$ according to Eq. 1. Hence, for any measurable $U\subseteq\mathcal{U}$, we have
@@
where $\nu_{0\mid t}^{{\boldsymbol{a}}_{t}}$ is the conditional measure of the reverse diffusion process given a noisy ${\boldsymbol{a}}_{t}$. Denote by $\hat{{\boldsymbol{a}}}_{0}({\boldsymbol{a}}_{t})=\mathbb{E}[{\boldsymbol{a}}_{0}\mid{\boldsymbol{a}}_{t}]$ the expected value of this measure. We will sometimes write $\hat{{\boldsymbol{a}}}_{0}$ to simplify the notation when the dependency on ${\boldsymbol{a}}_{t}$ is clear from context. Approximating $\nu_{0\mid t}^{{\boldsymbol{a}}_{t}}\approx\delta[\hat{{\boldsymbol{a}}}_{0}]$ with its mean, under the assumption that ${\boldsymbol{A}}({\boldsymbol{a}}_{0})\in H(\eta)$ is an element of the Cameron-Martin space of $\eta$,
@@
Note further that Eq. 7 shows that if the prior $\nu_{0}$ is such that ${\boldsymbol{A}}({\boldsymbol{a}}_{0})\in H(\eta)$ almost surely, then $\mu_{t}^{{\boldsymbol{a}}_{t}}\ll\eta$. In all, we have shown
$\frac{\,\mathrm{d}\mu_{t}^{{\boldsymbol{a}}_{t}}}{\,\mathrm{d}\eta}({\boldsymbol{u}})\approx\frac{\,\mathrm{d}\eta^{{\boldsymbol{A}}(\hat{{\boldsymbol{a}}}_{0})}}{\,\mathrm{d}\eta}({\boldsymbol{u}}),$ (9)
so that carrying this approximation over into the log-likelihood, we obtain $\tilde{\Phi}_{t}({\boldsymbol{a}}_{t},{\boldsymbol{u}})\approx\Phi(\hat{{\boldsymbol{a}}}_{0},{\boldsymbol{u}})$ as defined by Eq. 3. Thus, when we are able to calculate $\nabla_{{\boldsymbol{a}}_{t}}\Phi(\hat{{\boldsymbol{a}}_{0}}({\boldsymbol{a}}_{t}),{\boldsymbol{u}})$, we may substitute this expression into Eq. 6 in order to guide our unconditional diffusion process to approximately sample from the posterior $\nu^{{\boldsymbol{u}}}$.
@@
$\Phi(\hat{{\boldsymbol{a}}}_{0},{\boldsymbol{u}})=-\tfrac{1}{2}\left(\lVert{\boldsymbol{u}}-{\boldsymbol{A}}\hat{{\boldsymbol{a}}}_{0}\rVert^{2}_{H(\eta)}-\lVert{\boldsymbol{u}}\rVert^{2}_{H(\eta)}\right)$ (10)
@@
as the two differ by a constant depending only on ${\boldsymbol{u}}$. However, when $\mathcal{U}$ is infinite dimensional, the Cameron-Martin space $H(\eta)$ has measure zero under $\eta$, and so ${\boldsymbol{u}}-{\boldsymbol{A}}(\hat{{\boldsymbol{a}}_{0}})$ almost surely not be an element of $H(\eta)$ in which case Eq. 10 is ill-defined.

### v2

@@ Abstract
We propose a general framework for conditional sampling in PDE-based inverse problems, targeting the recovery of whole solutions from extremely sparse or noisy measurements. This is accomplished by a function-space diffusion model and plug-and-play guidance for conditioning. Our method first trains an unconditional discretization-agnostic denoising model using neural operator architectures. At inference, we refine the samples to satisfy sparse observation data via a gradient-based guidance mechanism. Through rigorous mathematical analysis, we extend Tweedie’s formula to infinite-dimensional Banach spaces, providing the theoretical foundation for our posterior sampling approach. Our method (FunDPS) accurately captures posterior distribution in function spaces under minimal supervision and severe data scarcity. Across five PDE tasks with only 3% observation, our method achieves an average 32% accuracy improvement over state-of-the-art fixed-resolution diffusion baselines while reducing sampling steps by 4x. Furthermore, multi-resolution fine-tuning ensures strong cross-resolution generalizability and speedup. To the best of our knowledge, this is the first diffusion-based framework to operate independently of discretization, offering a practical and flexible solution for forward and inverse problems in the context of PDEs. Code is available at https://github.com/neuraloperator/FunDPS.
@@ 1 Introduction
2) Multi-resolution conditional generation pipeline in function space (FunDPS).
We propose the function-space diffusion posterior sampling method that builds on the above theoretical framework, the pipeline of which is shown in Figure˜2. An unconditional diffusion model is first pretrained using the score matching objective as defined in [14]. The model works on a joint functional representation of PDE parameters and solutions, which allows us to solve forward problems, inverse problems, and a combination thereof using partial observations. The model employs a multi-resolution training strategy, which begins training on a coarse grid, then fine-tunes on a finer one. It reduces GPU training hours by 25%, which is enabled by the Gaussian random field (GRF)-based noise model and neural operator architecture. After training the unconditional model, we generate conditional samples as follows: The sample is initialized with GRF noise, followed by iterative denoising via a second-order deterministic sampler [16]. At each step, the sample evolves according to the conditional score operator, which decomposes into data prior (from the pretrained model) and observation-based likelihoods. Our generalized Tweedie’s formula enables us to efficiently approximate the measurement log-likelihood. FunDPS can also employ multi-resolution inference by performing most sampling steps at a lower resolution and only upsampling towards the end, which alone yields 2x speedup while maintaining accuracy.
@@
[…]
@@ 3.1 Problem Settings
We assume that the measurement noise $\varepsilon$ is a Gaussian random field (GRF) with covariance operator $\mathbf{C}_{\eta}$, i.e., $\varepsilon$ is drawn from the Gaussian measure $\eta=\mathcal{N}(0,\mathbf{C}_{\eta})$ on $\mathcal{U}$, which is the natural extension of Gaussian noise to function spaces [50]. Observe that, given $\boldsymbol{a}$}, we have that ${\boldsymbol{u}}\mid{\boldsymbol{a}}$ $\sim\eta^{{\boldsymbol{A}}({\boldsymbol{a}})}=\mathcal{N}({\boldsymbol{A}}({\boldsymbol{a}}),\mathbf{C}_{\eta})$ is drawn from a Gaussian measure $\eta^{{\boldsymbol{A}}({\boldsymbol{a}})}$ with mean ${\boldsymbol{A}}({\boldsymbol{a}})$. When the noise-free observation ${\boldsymbol{A}}({\boldsymbol{a}})$ is an element of the Cameron-Martin space $\mathbf{C}_{\eta}^{1/2}(\mathcal{U})=:H(\eta)\subset\mathcal{U}$ associated with $\mathbf{C}_{\eta}$, Cameron-Martin theorem allows us to compute the Radon-Nikodym derivative  [^2] We recall that the inner product and norm on the Cameron-Martin space $\mathcal{U}_{0}$ are readily computed via $\langle{\boldsymbol{u}}_{0},{\boldsymbol{u}}_{1}\rangle_{H(\eta)}=\langle\mathbf{C}_{\eta}^{-1/2}{\boldsymbol{u}}_{0},\mathbf{C}_{\eta}^{-1/2}{\boldsymbol{u}}_{1}\rangle_{\mathcal{U}},\qquad{\boldsymbol{u}}_{0},{\boldsymbol{u}}_{1}\in H(\eta).$
[…]
In other words, Eq.˜3 gives the density of $\eta^{\boldsymbol{a}}$ with respect to the noise measure $\eta$. We will use $\Phi:\mathcal{A}\times\mathcal{U}\to\mathbb{R}$ to represent the function $\Phi({\boldsymbol{a}},{\boldsymbol{u}})=\langle{\boldsymbol{u}},{\boldsymbol{A}}({\boldsymbol{a}})\rangle_{\mathcal{U}_{0}}-\tfrac{1}{2}\lVert{\boldsymbol{A}}({\boldsymbol{a}})\rVert^{2}_{\mathcal{U}_{0}}$. Loosely speaking, $\Phi({\boldsymbol{a}},{\boldsymbol{u}})$ is the log-likelihood of $\boldsymbol{u}$} given $\boldsymbol{a}$}. In the special case that ${\boldsymbol{u}}\in\mathbb{R}^{n}$ is finite dimensional (for instance, when $\boldsymbol{A}$} is the composition of a PDE solution operator and a finite observation mask), the measure $\eta$ corresponds to a mean-zero Gaussian random variable with covariance matrix $\mathbf{C}_{\eta}\in\mathbb{R}^{n\times n}$ . In this case, when $\mathbf{C}_{\eta}$ has full rank, the Cameron-Martin space is $\mathbf{C}_{\eta}^{1/2}(\mathbb{R}^{n})=\mathbb{R}^{n}$ so that the log-likelihood is defined for any value of ${\boldsymbol{A}}({\boldsymbol{a}})$.
@@
We propose to learn the prior distribution $\nu({\boldsymbol{a}})$ from data using the recent extension of diffusion models to function spaces [14, 51]. These methods define a sequence of distributions $\nu_{t}({\boldsymbol{a}}_{t})$ that progressively add noise to the data until approximately reaching a tractable latent distribution $\mu_{T}\approx\mathbf{\Gamma}$. Learning a score approximator ${\boldsymbol{D}}_{\theta}({\boldsymbol{a}}_{t},t)\approx\mathbb{E}[{\boldsymbol{a}}_{0}|{\boldsymbol{a}}_{t}]$ from the data via a variant of score matching, one can approximately reverse the noising process [^3] Following common conventions, we will also refer to the noising process as forward process. and sample from the data distribution. Moreover, by replacing the score with a conditional expectation, one can solve the reverse SDE given initial conditions [43]. Notably, the denoiser ${\boldsymbol{D}}_{\theta}$ is parametrized as a neural operator [14, 52], and the noise is chosen to be Gaussian random fields, as opposed to neural networks and multivariate Gaussian random variables for the finite-dimensional setting (see Appendix˜A for more details). The diffusion prior will be updated by (approximated) conditional likelihood during inference in order to solve inverse problems.
@@ 3.2 Conditional Score via Likelihood and Prior Score
for some specified time-dependent constants $\sigma_{t}>0$ [14]. Here, $\gamma=\mathcal{N}(0,\mathbf{C}_{\gamma})$ is now a Gaussian measure on the space $\mathcal{A}$. We will also use $\gamma_{t}=\mathcal{N}(0,\sigma_{t}^{2}\mathbf{C}_{\gamma})$ to represent the corresponding Gaussian measure with scaled covariance. The spaces $H(\gamma),H(\gamma_{t})$ represent the Cameron-Martin spaces associated with these Gaussian measures. We will write $\nu_{0}({\boldsymbol{a}}_{0})=\nu({\boldsymbol{a}})$ for the prior distribution over $\boldsymbol{a}$} and $\nu_{t}({\boldsymbol{a}}_{t})$ for the marginal distribution of ${\boldsymbol{a}}_{t}$ obtained during the forward process. Similarly, the conditional forward process is obtained in the same manner as Eq.˜4 except with initial conditions ${\boldsymbol{a}}_{0}\sim\nu^{{\boldsymbol{u}}}$ drawn from the posterior corresponding to a given, fixed $\boldsymbol{u}$}. We will use $\nu_{0}^{{\boldsymbol{u}}}=\nu^{{\boldsymbol{u}}}$ and $\nu_{t}^{{\boldsymbol{u}}}$ for the corresponding measures. Informally, $\nu_{t}^{{\boldsymbol{u}}}$ can be thought of as $\nabla_{{\boldsymbol{a}}}p_{t}({\boldsymbol{a}}_{t}|u)$.
As in prior work [14, 43], we make the assumption that the prior $\nu_{0}$ is supported on $H(\gamma)$, i.e., $\nu(H(\gamma))=1$. Note that as the posterior $\nu_{0}^{{\boldsymbol{u}}}\ll\nu_{0}$ is absolutely continuous with respect to the prior, we also have that $\nu_{0}^{{\boldsymbol{u}}}(H(\gamma))=1$. Under this assumption, both $\nu_{t}$ and $\nu_{t}^{{\boldsymbol{u}}}$ are equivalent to $\gamma_{t}$ in the sense of mutual absolute continuity (see Appendix˜C and [14, Lemma 13]).
@@ 3.3 Approximating the Likelihood in Function Spaces
While the calculations in the previous section in principle allow for plug-and-play posterior sampling, a major difficulty is that the time-dependent likelihood $\tilde{\Phi}_{t}$ is intractable. In particular, from Eq.˜3 we know the log-likelihood of $\boldsymbol{u}$} given a noise-free sample ${\boldsymbol{a}}_{0}$, but Eq.˜6 requires us to have access to the log-likelihood of $\boldsymbol{u}$} given a noisy sample ${\boldsymbol{a}}_{t}$. In the next two sections, we develop a method for approximating this likelihood through a conditional expectation.
Let us write $\mu_{t}^{{\boldsymbol{a}}_{t}}$ for the conditional distribution of ${\boldsymbol{u}}\mid{\boldsymbol{a}}_{t}$. This measure can be sampled from by first predicting the clean ${\boldsymbol{a}}_{0}$ from the noisy ${\boldsymbol{a}}_{t}$, followed by sampling from the noise measure $\eta^{{\boldsymbol{A}}({\boldsymbol{a}}_{0})}$ according to Eq.˜1. Hence, for any measurable $U\subseteq\mathcal{U}$, we have
@@
where $\nu_{0\mid t}^{{\boldsymbol{a}}_{t}}$ is the conditional measure of the reverse diffusion process given a noisy ${\boldsymbol{a}}_{t}$. Denote by $\hat{{\boldsymbol{a}}}_{0}({\boldsymbol{a}}_{t})=\mathbb{E}[{\boldsymbol{a}}_{0}\mid{\boldsymbol{a}}_{t}]$ the expected value of this measure. (We will sometimes write $\hat{{\boldsymbol{a}}}_{0}$ to simplify the notation when the dependency on ${\boldsymbol{a}}_{t}$ is clear from context.) Approximating $\nu_{0\mid t}^{{\boldsymbol{a}}_{t}}\approx\delta[\hat{{\boldsymbol{a}}}_{0}]$ with its mean, under the assumption that ${\boldsymbol{A}}({\boldsymbol{a}}_{0})\in H(\eta)$ is an element of the CM space of $\eta$, we have
@@
Note further that Eq.˜7 shows that if the prior $\nu_{0}$ is such that ${\boldsymbol{A}}({\boldsymbol{a}}_{0})\in H(\eta)$ almost surely, then $\mu_{t}^{{\boldsymbol{a}}_{t}}\ll\eta$. In all, we have shown
$\frac{\,\mathrm{d}\mu_{t}^{{\boldsymbol{a}}_{t}}}{\,\mathrm{d}\eta}({\boldsymbol{u}})\approx\frac{\,\mathrm{d}\eta^{{\boldsymbol{A}}(\hat{{\boldsymbol{a}}}_{0})}}{\,\mathrm{d}\eta}({\boldsymbol{u}}).$ (9)
Carrying this approximation over into the log-likelihood, we obtain $\tilde{\Phi}_{t}({\boldsymbol{a}}_{t},{\boldsymbol{u}})\approx\Phi(\hat{{\boldsymbol{a}}}_{0},{\boldsymbol{u}})$ as defined by Eq.˜3. Thus, when we are able to calculate $\nabla_{{\boldsymbol{a}}_{t}}\Phi(\hat{{\boldsymbol{a}}_{0}}({\boldsymbol{a}}_{t}),{\boldsymbol{u}})$, we may substitute this expression into Eq.˜6 in order to guide the diffusion process to approximately sample from the posterior $\nu^{{\boldsymbol{u}}}$.
@@
[…]
@@
[…]